OpenStudy (sushi121212):

Find the zeros of the function: f(x) x^2-4x√5 +19

satellite73 (satellite73):

guess you can use the quadratic formula

satellite73 (satellite73):

it is really \[f(x)=x^2-4\sqrt{5}x+19\]?

OpenStudy (sushi121212):

It says \[f(x)=x^2-4x \sqrt{5}+19\]

satellite73 (satellite73):

same thing

satellite73 (satellite73):

use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \[a=1, b=-4\sqrt5, c=19\]

OpenStudy (sushi121212):

I got \[\frac{ 4\sqrt{5}\pm2 }{ 2 }\]

OpenStudy (sushi121212):

Do I have to make it \[2\sqrt{5}\pm1\] instead?

zepdrix (zepdrix):

Yes, good job!